# Adversarial Disentanglement by Backpropagation with Physics-Informed Variational Autoencoder

### v2

@@ 5.2 Oscillator case study
(a) Mean and ±$\pm 2\sigma$ uncertainty bounds of the reconstructed response without adversarial training ($\lambda=-1$).
(b) Mean and ±$\pm 2\sigma$ uncertainty bounds of the reconstructed response with adversarial training ($\lambda=1/128$).
@@
Figure 9: $R^{2}$ value per subset of the latent variables and generative factor as a function of $\lambda$, averaged over $6$ runs. The shaded intervals correspond to two standard deviations.
It can be seen that the sign of $\lambda$ determines the nature of the training procedure, with positive values resulting in adversarial training. For negative values of $\lambda$ the training becomes collaborative, in the sense that the encoder attempts to find approximate posterior distributions over $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that are jointly informative about their respective modality as well as the response measurements $\bm{x}$. The magnitude of $\lambda$ determines the strength of the adversarial or collaborative training. To aid in the interpretation of the results, the behavior of the model is classified into four regimes. When $\lambda$ approaches $-1$ from above, the training is strongly collaborative, and the tasks of minimizing the error in the reconstruction of $\bm{x}$ and the prediction of the domain and class variables $\bm{c}$ and $\bm{y}$ are jointly prioritized. This is reflected by the relatively high scores obtained by the subsets $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ for the generative factors $m$ and $x_{\mathrm{0}}$. For $\lambda\rightarrow 0^{-}$, the auxiliary tasks are prioritized over the main task, and the amount of information about $m$ and $x_{\mathrm{0}}$ that is encoded in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ is limited. In this regime the behavior can be characterized as weakly collaborative. Conversely, for small positive values of $\lambda$ the training becomes weakly adversarial. In this regime the encoder will seek latent codes over $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that are uninformative about $\bm{x}$. Further increasing the GRL coefficient such that $\lambda\rightarrow 1$ yields a strongly adversarial model, and any information that can be used to reconstruct the domain and class variables is heavily weighted against the potential improvement in the reconstruction of $\bm{x}$. In this case the encoder fails to capture the variation in any of the generative factors.
@@ 5.3 Bridge case study
The final synthetic case study utilizes the two-span bridge benchmark presented in [Tatsis2019], illustrated in Figure 1(c). Members of a homogeneous population [Bull2021] of bridges are subjected to controlled loading tests, where a vehicle with known mass and moving at a constant velocity is used to excite the bridge response. The response is obtained as a strain influence line, expressed in parts per thousand (‰), measured by a point strain gauge placed at a distance of $5.625$ m from the start of the bridge, and at a height of $0.1$ m from the bottom of the cross section. Each time-series is composed of $64$ measurements, equally spaced in time $t\in[1,21]$ s, where $t=0$ s is the moment the vehicle enters the bridge. The behavior of each bridge is partially determined by the unknown vertical stiffnesses of the supports $k_{\mathrm{v},1}$, $k_{\mathrm{v},2}$ and $k_{\mathrm{v},3}$, which are taken to vary between different bridges due to variability in the design, construction and soil conditions. The boundary conditions are known to be symmetric such that $k_{\mathrm{v,1}}=k_{\mathrm{v,3}}$. The base-10 logarithms of the vertical stiffnesses are considered as physics-grounded latent variables. In the horizontal direction, only the left support has a large stiffness, while the rest are unconstrained. It is assumed that the position of the central pier can vary between members of the population by up to ±$\pm 1.0$ m from $L/2$. The nominal physics-based model includes a parameter $\delta_{\mathrm{s}}$, describing the offset of the pier relative to the center of the bridge in the longitudinal direction. Furthermore, fluctuations from the prescribed reference vehicle velocity $v_{\mathrm{ref}}=1$ m/s were observed during the tests that can not be accounted for in the nominal physics-based model. These fluctuations are modeled as a multiplicative term $\delta_{\mathrm{v}}$ such that $v=\delta_{\mathrm{v}}\cdot v_{\mathrm{ref}}$. Noisy measurements of the vehicle velocity, and the known pier offsets $\delta_{\mathrm{s}}$ are included as domain variables. It is assumed that the bridge decks are prone to deterioration in a region around the supports. During inspections performed by experts, each bridge is assigned scores $\bm{y}=(y_{1},y_{2})$, quantifying the deterioration of the structure near the left and middle supports respectively. Each $y_{i}$ takes values between zero and one, where zero represents pristine condition and unity corresponds to severe damage of the cross-section at that position. These scores are considered as class observables. Finally, a small variability $F_{\mathrm{ref}}$ is considered in the vehicle load such that $F=\delta_{\mathrm{F}}\cdot F_{\mathrm{ref}}$, where $F_{\mathrm{ref}}=100$ kN is the reference load. This variability is caused by deviation in the transverse position of the vehicle, and is considered as an unknown confounding influence. The response, domain and class observables are contaminated with i.i.d. samples of Gaussian white noise with standard deviations $\sigma_{\mathrm{x}}=\sigma_{\mathrm{c}}=\sigma_{\mathrm{y}}=10^{-4}$. The quantities involved in the case study along with their prior and ground truth distributions are summarized in Table 3. It is important to note that this case study is not intended to be a realistic representation of system identification and SHM for bridges, since it circumvents several important practical difficulties such as ensuring the consistency and alignment of data collected over long time-scales from a large number of structures. Furthermore, it is assumed for simplicity that the degradation condition of the bridges does not change significantly within the amount of time required to obtain the dataset.
@@
$y_{1}$ Class - $\mathcal{U}(0.0,1.0)$ 0.1
$y_{2}$ Class - $\mathcal{U}(0.0,1.0)$ 0.1
$\delta_{\mathrm{v}}$ Domain - $\mathcal{U}(0.9,1.1)$ 1.0
@@
$\delta_{\mathrm{F}}$ Unknown - $\mathcal{U}(0.9,1.1)$ 1.0
Training data for the full and nominal physical models is generated using the FE model of [Tatsis2019] as a simulator. The FE model is composed of quadrilateral isoparametric plane stress elements with $9$ nodes each, arranged in a $200\times 6$ grid. The length, width and height are assumed constant and equal to $L=25.0$ m, $w=0.1$ m and $h=0.6$ m for all of the bridges. The material is linear elastic, with Young’s modulus $E=200$ GPa, density $\rho=7850$ kg/m3 and Poisson’s ratio $\nu=0.3$. All supports are modeled as linear springs in the vertical direction. The equations of motion are integrated from $t_{0}=0$ seconds to $t_{1}=25$ seconds with a timestep of $\mathrm{d}t=0.00025$ seconds, using an implicit Newmark scheme with parameters $\gamma=1/2$ and $\beta=1/6$. The deterioration is modeled as a reduction of the cross section width, ranging from $0\%$ (for $y_{i}=0.0$) to $90\%$ (for $y_{i}=1.0$) in the region spanning $L/20$ around the corresponding support. It is noted that the deterioration is intentionally exaggerated to a large extent to ensure that it can be observed in the strain influence lines. The damaged regions are illustrated in Figure 1(c). The nominal physics-based model is assumed to be a simplistic but representative model for the behavior of the bridges in the population under study in their nominal condition, obtained through the procedure described in Section 5. Given the vertical stiffness of the abutments and support, and the offset of the central pier, the nominal model returns a time-series of strains.
The model is trained with $\lambda=1/1024$ and $d_{z_{\mathrm{c}}}=d_{z_{\mathrm{y}}}=4$. The predictions generated by the model while traversing each of the generative factors are shown in Figure 10. It can be seen that the data-driven component of the decoder is prevented from capturing variability in the reconstructed response when varying $\log_{10}k_{\mathrm{v},1}$, $\log_{10}k_{\mathrm{v},2}$ and $\delta_{\mathrm{F}}$, but is able to contribute to the components caused by the variation of the domain and class generative factors. Furthermore, the figure illustrates that the unknown confounder $\delta_{\mathrm{F}}$ can be partially accounted for by the physics based model. This is in contrast to the oscillator example (Section 5.2) where the influence of the unknown confounder could not be accounted for by the known physics.
@@
Figure 10: Mean prediction and ±$\pm 2\sigma$ uncertainty bounds for the physics-based $\hat{\bm{x}}_{\mathrm{p}}$ and data-driven $\hat{\bm{x}}_{\mathrm{d}}$ components, and combined prediction $\hat{\bm{x}}$ while traversing the generative factors. The input response measurements are denoted as dots in the bottom row.
The previous conclusions are further supported by the traversal of the latent space, shown in Figure 11, which indicates that the domain and class subsets of the latent variables encode information that enables the auxiliary decoders to predict the domain and class labels, and the response decoder to correct the physics-based model prediction. It can be seen that the influence of the unknown confounder $\delta_{\mathrm{F}}$ is partly captured as variability in the physics-based subset $\bm{z}_{\mathrm{x}}$, indicating that model form uncertainty is compensated by inferring an “effective” value of the physics-grounded latent variables. Figure 11 also suggests that the domain and class latent variables only capture variability in the corresponding generative factors, whereas the physics-based latent variables are always active, providing additional evidence for the claim that the adversarial objective induces a disentangled representation while prioritizing the use of the known physics. Importantly, Figure 10 and Figure 11 illustrate that the adversarial training can feasibly constrain $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$, such that it only contributes to the prediction when justified by additional domain and class observables.
@@
As discussed in Section 3, the model is trained in a fully supervised manner to simultaneously reconstruct the domain and class variables from the input measurements, making it possible to handle tasks such as damage detection, where predicting the class labels $\bm{y}$ from input measurements $\bm{x}$ is of interest. Given a trained model, the condition labels of a similar bridge can be predicted from response measurements. The performance is evaluated in two different cases, illustrated in Figure 12, referred to as “interpolation” and “extrapolation” respectively. For each case, the space of physics-based generative factors is subdivided into four quarters. In the interpolation case, the model is trained on $N_{\mathrm{train}}=1024$ samples from three quarters and evaluated on $N_{\mathrm{test}}=512$ samples from the fourth. In the extrapolation case, the model is trained on data from a single quarter and evaluated on the remaining three, using the same train and test set sizes. All other generative factors are sampled from the ground truth distributions presented in Table 3. To obtain a more comprehensive evaluation, each of the two cases is divided into four sub-cases, over which the results are averaged.
@@
The proposed model using two different hyperparameter settings, denoted as DPIVAE-A and DPIVAE-B, is compared with linear regression (LIN), Gaussian Process Regression (GPR) and a Multi-Layer Perceptron (MLP). For DPIVAE-A the GRL is not utilized, i.e. $\lambda=-1$, and separate encoders are used for each subset of the latent variables. For DPIVAE-B the GRL hyperparameter is taken as $\lambda=1/1024$. The GPR is implemented with a Radial Basis Function kernel and additive Gaussian white noise. The MLP is formulated with two hidden layers, each with a width of $64$ units and a Rectified Linear Unit (ReLU) activation function. Results in terms of interpolation and extrapolation performance of each model is quantified in terms of the R2 and Mean Squared Error (MSE), shown in Table 4. These results are intended to highlight that the performance of the different models is comparable, and that the proposed model can feasibly be used to predict the class variables in a complex high-dimensional case study. Manual hyperparameter tuning is performed for the models involved in the comparison.
Table 4: Mean and standard deviation of $R^{2}$ and $\mathrm{MSE}$ for the task of predicting $\bm{y}$, averaged over $6$ runs.
@@
Model R2()($\uparrow$)( ↑ ) MSE()($\downarrow$)( ↓ ) R2()($\uparrow$)( ↑ ) MSE()($\downarrow$)( ↓ )
@@ 6 Discussion
The results presented in Section 5 indicate that the proposed architecture and adversarial objective effectively constrain the posterior distribution over domain and class latent variables, and by extension, the flexibility of the data-driven decoder components. The strength of the regularization is controlled by an interpretable hyperparameter that determines the strength of the gradient reversal. This allows for the main and auxiliary decoders to be trained in a collaborative or adversarial manner. This hyperparameter effectively controls the relative importance of the physics-based and data-driven components, and can be used to encourage the model to preferentially utilize the known physics. When the training is adversarial, the domain and class latent spaces encode features of the response measurements that can be related to the domain or class variables, and that can not be accounted for by the known physics, and are simultaneously constrained to avoid overriding the physics-based model predictions. Because neither the domain or class observables are necessary during model evaluation, the proposed approach has the potential to reduce the need for cumbersome and expensive data collection methods, such as those involving elaborate experimental procedures or expert assessments.

### v3

@@ 5.2.2 Model behavior in the presence of unknown confounders
(a) Mean and $\pm 2\sigma$ uncertainty bounds of the reconstructed response without adversarial training ($\lambda=-1$).
(b) Mean and $\pm 2\sigma$ uncertainty bounds of the reconstructed response with adversarial training ($\lambda=1/128$).
@@ 5.2.3 Quantitative assessment of disentanglement
Figure 9: $R^{2}$ value per subset of the latent variables and generative factor as a function of $\lambda$, averaged over 66 runs. The shaded intervals correspond to two standard deviations.
It can be seen that the sign of $\lambda$ determines the nature of the training procedure, with positive values resulting in adversarial training. For negative values of $\lambda$ the training becomes collaborative, in the sense that the encoder attempts to find approximate posterior distributions over $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that are jointly informative about their respective modality as well as the response measurements $\bm{x}$. The magnitude of $\lambda$ determines the strength of the adversarial or collaborative training. To aid in the interpretation of the results, the behavior of the model is classified into four regimes. When $\lambda$ approaches 1-1 from above, the training is strongly collaborative, and the tasks of minimizing the error in the reconstruction of $\bm{x}$ and the prediction of the domain and class variables $\bm{c}$ and $\bm{y}$ are jointly prioritized. This is reflected by the relatively high scores obtained by the subsets $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ for the generative factors $m$ and $x_{\mathrm{0}}$. For $\lambda\rightarrow 0^{-}$, the auxiliary tasks are prioritized over the main task, and the amount of information about $m$ and $x_{\mathrm{0}}$ that is encoded in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ is limited. In this regime the behavior can be characterized as weakly collaborative. Conversely, for small positive values of $\lambda$ the training becomes weakly adversarial. In this regime the encoder will seek latent codes over $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that are uninformative about $\bm{x}$. Further increasing the GRL coefficient such that $\lambda\rightarrow 1$ yields a strongly adversarial model, and any information that can be used to reconstruct the domain and class variables is heavily weighted against the potential improvement in the reconstruction of $\bm{x}$. In this case the encoder fails to capture the variation in any of the generative factors.
@@ 5.3 Bridge case study
5.3.1 Case study description
The final synthetic case study utilizes the two-span bridge benchmark presented in [64], illustrated in Figure 1(c). Members of a homogeneous population [9] of bridges are subjected to controlled loading tests, where a vehicle with known mass and moving at a constant velocity is used to excite the bridge response. The response is obtained as a strain influence line, expressed in parts per thousand (‰), measured by a point strain gauge placed at a distance of $5.625$ m from the start of the bridge, and at a height of $0.1$ m from the bottom of the cross section. Each time-series is composed of $64$ measurements, equally spaced in time $t\in[1,21]$ s, where $t=0$ s is the moment the vehicle enters the bridge.
@@ 5.3.1 Case study description
The behavior of each bridge is partially determined by the unknown vertical stiffnesses of the supports $k_{\mathrm{v},1}$, $k_{\mathrm{v},2}$ and $k_{\mathrm{v},3}$, which are taken to vary between different bridges due to variability in the design, construction and soil conditions. The boundary conditions are known to be symmetric such that $k_{\mathrm{v,1}}=k_{\mathrm{v,3}}$. The base-10 logarithms of the vertical stiffnesses are considered as physics-grounded latent variables. In the horizontal direction, only the left support has a large stiffness, while the rest are unconstrained. It is assumed that the position of the central pier can vary between members of the population by up to $\pm 1.0$ m from $L/2$. Furthermore, fluctuations from the prescribed reference vehicle velocity $v_{\mathrm{ref}}=1$ m/s were observed during the tests that can not be accounted for in the nominal physics-based model. These fluctuations are modeled as a multiplicative term $\delta_{\mathrm{v}}$ such that $v=\delta_{\mathrm{v}}\cdot v_{\mathrm{ref}}$. Noisy measurements of the vehicle velocity, and the known pier offsets $\delta_{\mathrm{s}}$ are included as domain variables. It is assumed that the bridge decks are prone to deterioration in a region around the supports. During inspections performed by experts, each bridge is assigned scores $\bm{y}=(y_{1},y_{2})$, quantifying the deterioration of the structure near the left and middle supports respectively. Each $y_{i}$ takes values between zero and one, where zero represents pristine condition and unity corresponds to severe damage of the cross-section at that position. These scores are considered as class observables. Finally, a small variability is considered in the vehicle load such that $F=\delta_{\mathrm{F}}\cdot F_{\mathrm{ref}}$, where $F_{\mathrm{ref}}=100$ kN is the reference load. This variability is caused by deviation in the transverse position of the vehicle, and is considered as an unknown confounding influence. The quantities involved in the case study along with their prior and ground truth distributions are summarized in Table 3
@@
$y_{1}$ - Class - $\mathcal{U}(0.0,1.0)$ 0.1
$y_{2}$ - Class - $\mathcal{U}(0.0,1.0)$ 0.1
$\delta_{\mathrm{v}}$ - Domain - $\mathcal{U}(0.9,1.1)$ 1.0
@@
$\delta_{\mathrm{F}}$ - Unknown - $\mathcal{U}(0.9,1.1)$ 1.0
Training data for the full and nominal physical models is generated using the FE model of [64] as a simulator. The FE model is composed of quadrilateral isoparametric plane stress elements with $9$ nodes each, arranged in a $200\times 6$ grid. The length, width and height are assumed constant and equal to $L=25.0$ m, $w=0.1$ m and $h=0.6$ m for all of the bridges. The material is linear elastic, with Young’s modulus $E=200$ GPa, density $\rho=7850$ kg/m3 and Poisson’s ratio $\nu=0.3$. All supports are modeled as linear springs in the vertical direction. The equations of motion are integrated from $t_{0}=0$ seconds to $t_{1}=25$ seconds with a timestep of $\mathrm{d}t=0.00025$ seconds, using an implicit Newmark scheme with parameters $\gamma=1/2$ and $\beta=1/6$. The deterioration is modeled as a reduction of the cross section width, ranging from $0\%$ (for $y_{i}=0.0$) to $90\%$ (for $y_{i}=1.0$) in the region spanning $L/20$ around the corresponding support. It is noted that the deterioration is intentionally exaggerated to a large extent to ensure that it can be observed in the strain influence lines. The damaged regions are illustrated in Figure 1(c).
The nominal physics-based model is assumed to be a simplistic but representative model for the behavior of the bridges in the population under study in their nominal condition, obtained through the procedure described in Section 5. In addition to the physics-grounded latent variables, the nominal model also includes the parameter $\delta_{\mathrm{s}}$ as an input, describing the offset of the pier relative to the center of the bridge in the longitudinal direction. Given the vertical stiffness of the abutments and support and the offset of the central pier, the nominal model returns a time-series of strains.
The response, domain and class observables are contaminated with i.i.d. samples of Gaussian white noise with standard deviations $\sigma_{\mathrm{x}}=\sigma_{\mathrm{c}}=\sigma_{\mathrm{y}}=10^{-4}$. It is important to note that this case study is not intended to be a realistic representation of system identification and SHM for bridges, since it circumvents several important practical difficulties such as ensuring the consistency and alignment of data collected over long time-scales from a large number of structures. Furthermore, it is assumed for simplicity that the degradation condition of the bridges does not change significantly within the amount of time required to obtain the dataset.
5.3.2 Qualitative assessment of disentanglement
The model is trained with $\lambda=1/1024$ and $d_{z_{\mathrm{c}}}=d_{z_{\mathrm{y}}}=4$. The predictions generated by the model while traversing each of the generative factors are shown in Figure 10. It can be seen that the data-driven component of the decoder is prevented from capturing variability in the reconstructed response when varying $\log_{10}k_{\mathrm{v},1}$, $\log_{10}k_{\mathrm{v},2}$ and $\delta_{\mathrm{F}}$, but is able to contribute to the components caused by the variation of the domain and class generative factors. Furthermore, the figure illustrates that the unknown confounder $\delta_{\mathrm{F}}$ can be partially accounted for by the physics-based model. This is in contrast to the oscillator example (Section 5.2) where the influence of the unknown confounder could not be accounted for by the known physics.
@@ 5.3.2 Qualitative assessment of disentanglement
Figure 10: Mean prediction and $\pm 2\sigma$ uncertainty bounds for the physics-based $\hat{\bm{x}}_{\mathrm{p}}$ and data-driven $\hat{\bm{x}}_{\mathrm{d}}$ components, and combined prediction $\hat{\bm{x}}$ while traversing the generative factors. The input response measurements are denoted as dots in the bottom row.
The previous conclusions are further supported by the traversal of the latent space, shown in Figure 11, which indicates that the domain and class subsets of the latent variables encode information that enables the auxiliary decoders to predict the domain and class labels, and the response decoder to correct the physics-based model prediction. It can be seen that the influence of the unknown confounder $\delta_{\mathrm{F}}$ is partly captured as variability in the physics-based subset $\bm{z}_{\mathrm{x}}$, indicating that model form uncertainty is compensated by inferring an “effective” value of the physics-grounded latent variables. Figure 11 also suggests that the domain and class latent variables only capture variability in the corresponding generative factors, whereas the physics-grounded latent variables are always active, providing additional evidence for the claim that the adversarial objective induces a disentangled representation while prioritizing the use of the known physics. Importantly, Figure 10 and Figure 11 illustrate that the adversarial training can feasibly constrain $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$, such that it only contributes to the prediction when justified by additional domain and class observables.
@@
5.3.3 Application to damage identification
As discussed in Section 3, the model is trained in a fully supervised manner to simultaneously reconstruct the domain and class variables from the input measurements, making it possible to handle tasks such as damage detection, where predicting the class labels $\bm{y}$ from input measurements $\bm{x}$ is of interest. Given a trained model, the condition labels of a similar bridge can be predicted from response measurements. The performance is evaluated in two different cases, illustrated in Figure 12, referred to as “interpolation” and “extrapolation” respectively. For each case, the space of physics-grounded generative factors is subdivided into four quarters. In the interpolation case, the model is trained on $N_{\mathrm{train}}=1024$ samples from three quarters and evaluated on $N_{\mathrm{test}}=512$ samples from the fourth. In the extrapolation case, the model is trained on data from a single quarter and evaluated on the remaining three, using the same train and test set sizes. All other generative factors are sampled from the ground truth distributions presented in Table 3. To obtain a more comprehensive evaluation, each of the two cases is divided into four sub-cases, over which the results are averaged.
@@ 5.3.3 Application to damage identification
The proposed Disentangled Physics-Informed Variational Autoencoder (DPIVAE), using two different hyperparameter settings denoted as DPIVAE-A and DPIVAE-B, is compared with linear regression (LIN), Gaussian Process Regression (GPR) and a Multi-Layer Perceptron (MLP). For DPIVAE-A the GRL is not utilized, i.e. $\lambda=-1$, and separate encoders are used for each subset of the latent variables. For DPIVAE-B the GRL hyperparameter is taken as $\lambda=1/1024$. The GPR is implemented with a Radial Basis Function kernel and additive Gaussian white noise. The MLP is formulated with two hidden layers, each with a width of $64$ units and a Rectified Linear Unit (ReLU) activation function. Results in terms of interpolation and extrapolation performance of each model is quantified in terms of the R2 and Mean Squared Error (MSE), shown in Table 4. These results are intended to highlight that the performance of the different models is comparable, and that the proposed model can feasibly be used to predict the class variables in a complex high-dimensional case study. Manual hyperparameter tuning is performed for the models involved in the comparison.
Table 4: Mean and standard deviation of $R^{2}$ and $\mathrm{MSE}$ for the task of predicting $\bm{y}$, averaged over 66 runs.
@@
Model R2()($\uparrow$) MSE()($\downarrow$) R2()($\uparrow$) MSE()($\downarrow$)
@@ 6.1 Contributions and strengths
The results presented in Section 5 indicate that the proposed architecture and adversarial objective effectively constrain the posterior distribution over domain and class latent variables, and by extension, the flexibility of the data-driven decoder components. The constraint is controlled by an interpretable hyperparameter that determines the strength of the gradient reversal. This allows for the main and auxiliary decoders to be trained in a collaborative or adversarial manner. This hyperparameter effectively controls the relative importance of the physics-based and data-driven components, and can be used to encourage the model to preferentially utilize the known physics. When the training is adversarial, the domain and class latent spaces encode features of the response measurements that can be related to the observed domain or and class variables, and that can not be accounted for by the known physics. Simultaneously, the data-driven components of the model are constrained to avoid overriding the physics-based model predictions. Because neither the domain or class observables are necessary during model evaluation, the proposed approach has the potential to reduce the need for cumbersome and expensive data collection methods, such as those involving elaborate experimental procedures or expert assessments.
6.2 Assumptions and limitations